# The Use of Large Language Models (LLM) for Cyber Threat Intelligence (CTI) in Cybercrime Forums

### v1

@@ Abstract
0.5in0.5in Large language models (LLMs) can be used to analyze cyber threat intelligence (CTI) data from cybercrime forums, which contain extensive information and key discussions about emerging cyber threats. However, to date, the level of accuracy and efficiency of LLMs for such critical tasks has yet to be thoroughly evaluated. Hence, this study assesses the accuracy of an LLM system built on the OpenAI GPT-3.5-turbo model [7] to extract CTI information. To do so, a random sample of 500 daily conversations from three cybercrime forums—XSS, Exploit.in, and RAMP—was extracted, and the LLM system was instructed to summarize the conversations and code 10 key CTI variables, such as whether a large organization and/or a critical infrastructure is being targeted. Then, two coders reviewed each conversation and evaluated whether the information extracted by the LLM was accurate. The LLM system performed strikingly well, with an average accuracy score of 98%. Various ways to enhance the model were uncovered, such as the need to help the LLM distinguish between stories and past events, as well as being careful with verb tenses in prompts. Nevertheless, the results of this study highlight the efficiency and relevance of using LLMs for cyber threat intelligence.
@@ 1 Introduction
With the rise of large language models (LLMs), which are text-generating technologies trained on vast amounts of data [3, 4, 12], the scope and aim of artificial intelligence (AI) have changed. Now, AI can be used for a myriad of applications, such as writing stories on the fly or summarizing complex scientific content [11]. One key application where AI can be useful is cyber threat intelligence (CTI) [3, 4, 12]. Indeed, there exists a vast array of cybercrime conversations in forums on both the clear and dark web [1, 5]. These conversations often leak key information that can be used by companies and governments to detect and sometimes even prevent cyber attacks [1, 5]. The question is whether and how LLMs can accurately be used for cyber threat intelligence on such forums. Indeed, can we trust such technology for CTI? Can it replace first-level threat analysts, that is, analysts who read and extract relevant information from cybercrime forums?
This study assesses the extent to which an LLM system is accurate when extracting and summarizing information from cybercrime forums. Using a random sample of 500 daily conversations from three cybercrime forums—XSS, Exploit[.]in, and RAMP—we instructed an LLM to summarize the conversations and extract specific CTI information from them. CTI information included whether a sale was conducted, whether a large organization or a critical infrastructure was discussed, whether initial access to an organization was mentioned, whether a vulnerability that is remotely exploitable or actively exploited was mentioned, and whether users discussed geopolitical conflicts. Any technology mentioned in a conversation was also coded, as well as any industry, when available. Such information is key for CTI: it allows analysts to narrow down cybercrime conversations that focus, for example, on a specific technology in an industry or on specific vulnerabilities targeting large organizations.
To assess the accuracy of the LLM, two coders conducted a thorough review of each conversation and evaluated whether the information extracted by the LLM was accurate. This process 1) determined the level of accuracy of the technology, and 2) highlighted flaws that, once acknowledged, could be fixed.
In the end, the LLM system performed exceptionally well, accurately coding the ten variables, on average, 98% of the time with a minimum of 95% and a maximum of 100%. The assessment of the LLM also highlighted areas of improvement that could be useful for any researcher wishing to use LLMs to summarize forum conversations, including verb tenses in prompts and the impact of using large or vague concepts. The results highlight that large language models can be effectively used for cyber threat intelligence.
@@ 2.1.1 Data Collection and Preprocessing
The first step involves selecting and collecting data from high-quality cybercrime forums. These forums are continuously monitored, and new messages are extracted daily. The system processes both new and existing discussion threads, ensuring comprehensive coverage of relevant conversations.
@@ 2.1.2 Contextual Information Extraction
For each extracted message, the system determines the context based on whether the discussion thread is new or has been previously processed. If the thread is new, the title serves as the context. For existing threads, a summary of the prior conversation is used to provide context.
@@ 2.2 Key CTI Variables and Prompts
From daily cybercrime conversations, ten variables were extracted using specific prompts with the LLM. These variables were chosen because they provide key information for CTI, such as which technology or industry is targeted. Having readily accessible and easily identifiable information for CTI ensures that analysts can narrow down their focus to conversations relevant for their organization. Anyone wishing to reproduce the analysis could develop their own key variables. Table 1 presents the variables that were extracted by the LLM for this study.
@@
summary Generate concise extraction summaries with ample information for effective similarity searches, including conversation outcomes and all technical details that could be used by an analyst to investigate the threat.
is_sale An actor is selling something.
is_initial_access Involves the sale of initial access to corporate or organization network.
is_targeting_mainstream A mainstream software or hardware or a product utilized by many businesses and organizations for its features and security capabilities is being targeted or compromised.
is_targeting_large_organization A large organization is being targeted by the actors.
is_targeting_critical_infrastructure A critical infrastructure provider is targeted by the threat.
is_remotely_exploitable A mentioned vulnerability is remotely exploitable.
is_actively_exploitable A mentioned vulnerability is being actively exploited.
is_geopolitics The discussion involves geopolitical issues.
targeted_technologies Targeted or abused products or technology names.
industries Names of the industries relevant to the content, including those targeted by threat actors. Choose from the following options, each accompanied by a specific definition:
@@ 2.3 Coding Process
Once the individual coding was completed, the two analysts performed an inter-coder agreement to obtain a common decision for each unit summary. Such inter-coder agreement allowed them to pinpoint discrepancies, but also find areas of improvement to fine-tune the LLM system. In the end, the inter-coder agreement was high, with an average of 98.2%, a minimum of 96.2% and a maximum of 99.8%. A merged database was created resulting from the agreement. This merged database was used to determine the accuracy of the LLM system in coding each of the ten CTI variables from the 500 daily conversations on cybercrime forums. The results are presented below.
3 Assessing the accuracy of the LLM System for CTI
The LLM system performed strikingly well in coding variables that represented valuable CTI information from cybercrime forums. Indeed, on average, the LLM was accurate in 97.96% of cases, with a minimum accuracy of 95% for the variable is_targeting_critical_infrastructure and a maximum of 100% for the variable industries. The percentage accuracy for each variable is presented in Table 2 below.
@@ 3 Assessing the accuracy of the LLM System for CTI
summary 500 98.8%
is_sale 500 97.2%
is_initial_access 500 99.0%
is_targeting_large_organization 500 96.2%
is_targeting_critical_infrastructure 500 95.0%
is_remotely_exploitable 500 98.2%
is_actively_exploitable 500 99.4%
is_geopolitics 500 96.0%
targeted_technologies 115 99.1%
industries 273 100%
Table 2: LLM System Accuracy for each CTI Variable
For the summary variable, the coders ensured that the summary provided by the LLM system was accurate, and it was the case for 98.8% (N=500) of the conversation flows analyzed. However, it must be acknowledged that sometimes, the coders did notice that some key CTI information was missing from the summary. This is because “summarizing” a conversation means cutting text elements, which sometimes may be interpreted as important by some and not by others. The variable is_sale was also well coded by the LLM system, with an accuracy of 97.2%. Such variables captured when a user was conducting or announcing a sale. Most cases in which the LLM was wrong related to individuals expressing interest in sales or discussing sales, while not conducting or announcing any. The LLM system also performed really well when coding the variable is_initial_access, flagging messages involving the sale of initial access to a corporate or organizational network, with an accuracy of 99%. Hence, almost all posts that involve a sale and/or the sale of an initial access could be narrowed down using these two variables.
The variableis_targeting_large_organization had an accuracy of 96.2% and is_targeting
_critical_infrastructure
had the lowest accuracy of 95%. For these two variables, the LLM coding errors related mainly to definitions, specifically, defining what is a “large” organization or what is a “critical infrastructure”. Analysts could still use these two variables to assess which large organization and/or critical infrastructure is discussed in cybercrime forums.
The variable is_remotely_exploitable,flagged conversations in which a vulnerability that is remotely exploitable was mentioned. This variable was accurately coded 98.2% of the time. The LLM system also performed well (score of 99.4%) when coding the variable is_actively_exploitable, which aimed at flagging conversations in which a vulnerability actively exploited was mentioned in cybercrime conversations. These two variables combined can narrow down conversations on relevant vulnerabilities that organizations should prioritize in fixing.
The variable is_geopolitics aimed at flagging any conversations that related to geopolitical issues. It was accurately coded 96% of the time by the LLM. For this variable, the LLM tended to code the variable as TRUE when a country was mentioned, particularly if the country was associated with a tense geopolitical context. Hence, analysts could use this variable to narrow down conversations that may target specific countries or geopolitical conflicts.
@@
For the variable targeted_technologies, the coders made sure that the name of a technology mentioned in the message matched the technology indicated by the LLM. For all messages that had a technology (N=115), the accuracy was 99.1%, indicating that the LLM missed a technology really rarely. Given the high accuracy of the variable, analysts could use this variable to assess whether their technologies are discussed in cybercrime forums.
Finally, the variable industries was accurately coded 100% of the time, the highest accuracy score. This variable was coded when an industry was mentioned in a message (N=273). Such a high score might be due to this variable being well detailed, which seemed to contribute to the LLM’s effectiveness in correctly identifying the appropriate industry categories indicated in the messages. With this variable, analysts -and even policy makers- can narrow down conversations in cybercrime forums that target specific industries, thus facilitating their risks assessments.
@@ 4 Coders’ Insights to Fine-Tune the LLM System
Given the results of the analysis, there is no doubt that LLM systems can be useful in extracting key CTI information from cybercrime forums. Indeed, the summaries generated from daily conversations focused on relevant information, even when a large number of messages were posted. Such relevant information was sometimes even missed by the analysts. The LLM also demonstrated an exceptional ability to code key CTI variables, as evidenced by the results obtained: it had an accuracy of 98% on average. Such high results were unexpected and showcase interesting future research avenues.
Reviewing the same conversations allowed the coders to discuss and identify areas for improvement in coding the CTI variables. These areas are useful for any researcher or analyst wishing to use LLMs to code variables based on text input. They are presented below.
@@ 4.2 Verb Tenses in Prompts
Most likely, the verb tense in the prompt, ”an actor is selling something,” was the reason why the LLM did not code this conversation as involving a sale. However, upon reflection, it could be considered a sale because the item in question has been sold. In the end, if the goal is to code all sales, past or present, it might be necessary to revisit the verb tenses in the prompts given to the LLM.
@@ 4.5 Links between Variables with Similar Concepts
In this example, the LLM coded FALSE for is_targeting_critical_infrastructure and TRUE for is_targeting_large_organization. In the industries variable, the LLM indicated [’Critical Infrastructure’, ’Military’]. This discrepancy raises questions about the definitions of these three variables, as discussed above, and how the LLM understood them. It also showcases that some variables may be used to code others. For example, a technology mentioned in the summary was not always associated with an organization, despite the obvious connection between the two. In this excerpt:
@@ 4.7 The LLM System is Imperfect, just like Humans
The LLM coded the variable is_initial_access as FALSE even though the actor sells a tool that allows remote control over a target machine via a web browser. Hence, in some cases, the LLM missed the information. The LLM system was therefore not flawless, just like humans. Although rare, it made coding errors that could not be explained.

### v3

@@ Abstract
0.5in0.5in Large language models (LLMs) can be used to analyze cyber threat intelligence (CTI) data from cybercrime forums, which contain extensive information and key discussions about emerging cyber threats. However, to date, the level of accuracy and efficiency of LLMs for such critical tasks has yet to be thoroughly evaluated. Hence, this study assesses the performance of an LLM system built on the OpenAI GPT-3.5-turbo model [8] to extract CTI information. To do so, a random sample of more than 700 daily conversations from three cybercrime forums—XSS, Exploit.in, and RAMP—was extracted, and the LLM system was instructed to summarize the conversations and predict 10 key CTI variables, such as whether a large organization and/or a critical infrastructure is being targeted, with only simple human-language instructions. Then, two coders reviewed each conversation and evaluated whether the information extracted by the LLM was accurate. The LLM system performed well, with an average accuracy score of 96.23%, an average precision of 90% and an average recall of 88.2%. Various ways to enhance the model were uncovered, such as the need to help the LLM distinguish between stories and past events, as well as being careful with verb tenses in prompts. Nevertheless, the results of this study highlight the relevance of using LLMs for cyber threat intelligence.
@@ 1 Introduction
With the rise of large language models (LLMs), which are text-generating technologies trained on vast amounts of data [3, 4, 13], the scope and aim of artificial intelligence (AI) have changed. Now, AI can be used for a myriad of applications, such as writing stories on the fly or summarizing complex scientific content [12]. One key application where AI can be useful is cyber threat intelligence (CTI) [3, 4, 13]. Indeed, there exists a vast array of cybercrime conversations in forums on both the clear and dark web [1, 6]. These conversations often leak key information that can be used by companies and governments to detect and sometimes even prevent cyber attacks [1, 6]. The question is whether and how LLMs can accurately be used for cyber threat intelligence on such forums. Indeed, can we trust such technology for CTI? Can it replace first-level threat analysts, that is, analysts who read and extract relevant information from cybercrime forums? And at what cost?
This study assesses the extent to which an LLM system is accurate when extracting and summarizing information from cybercrime forums. Using a random sample of more than 700 daily conversations from three cybercrime forums—XSS, Exploit[.]in, and RAMP—we instructed an LLM to summarize the conversations and extract specific CTI information from them. CTI information included whether a sale was conducted, whether a large organization or a critical infrastructure was discussed, whether initial access to an organization was mentioned, whether a vulnerability that is remotely exploitable or actively exploited was mentioned, and whether users discussed geopolitical conflicts. Any technology mentioned in a conversation was also coded, as well as any industry, when available. Such information is key for CTI: it allows analysts to narrow down cybercrime conversations that focus, for example, on a specific technology in an industry or on specific vulnerabilities targeting large organizations.
It is worth mentioning as well that traditional machine learning classifiers meant to do that with extensive and expensive training data that were manually labelled and curated. The aim of this study is to assess the capacity of LLMs to perform these task without having to prepare such a training corpus, in a so-called zero-shot setting.
To assess the performance of the LLM, two coders conducted a thorough review of each conversation and evaluated whether the information extracted by the LLM was accurate. This process 1) determined the level of performance of the technology according to accuracy, precision and recall, and 2) highlighted flaws that, once acknowledged, could be fixed.
In the end, the LLM system performed well, with an average accuracy score of 96.23%, an average precision of 90% and an average recall of 88.2%. The assessment of the LLM also highlighted areas of improvement that could be useful for any researcher wishing to use LLMs to summarize forum conversations, including verb tenses in prompts and the impact of using large or vague concepts. The results validate that large language models can be effectively used for cyber threat intelligence.
@@ 2.1.1 Data Collection and Preprocessing
The first step involves selecting and collecting data from high-quality cybercrime forums. The system processes, every day, new daily conversations from new and existing threads on these cybercrime forums. These daily conversations (also called events) are then analyzed.
@@ 2.1.2 Contextual Information Extraction
For each new daily extractions, the system determines the context based on whether the discussion thread is new or has been previously processed. If the thread is new, the title serves as the context. For existing threads, a summary of the prior conversations is used to provide context.
@@ 2.2 Key CTI Variables and Prompts
From each event, ten variables were extracted using specific but simple prompts with the LLM. These variables were chosen because they provide key and actionable information for CTI, such as which technology or industry is targeted. Having readily accessible and easily identifiable information for CTI ensures that analysts can narrow down their focus to conversations relevant for their organization, or get alerted when matching certain criteria. Anyone wishing to reproduce the analysis could develop their own key variables. Table 1 presents the variables that were extracted by the LLM for this study and their related prompts.
@@
summary (char) Generate concise extraction summaries with ample information for effective similarity searches, including conversation outcomes and all technical details that could be used by an analyst to investigate the threat.
is_sale (1/0) An actor is selling something.
is_initial_access (1/0) Involves the sale of initial access to corporate or organization network.
is_targeting_large_organization (1/0) A large organization is being targeted by the actors.
is_targeting_critical_infrastructure (1/0) A critical infrastructure provider is targeted by the threat.
is_remotely_exploitable (1/0) A mentioned vulnerability is remotely exploitable.
is_actively_exploitable (1/0) A mentioned vulnerability is being actively exploited.
is_geopolitics (1/0) The discussion involves geopolitical issues.
targeted_technologies (list) Targeted or abused products or technology names.
industries (list) Names of the industries relevant to the content, including those targeted by threat actors. Choose from the following options, each accompanied by a specific definition:
@@ 2.3 Coding Process
The LLM model discussed above was then applied on these conversations. However, some variables had unbalanced results (i.e., had a small number of TRUE instances). To overcome this limit, we randomly selected additional conversations to make sure that each binary variable had at least 50 conversations scoring TRUE and 50 conversations scoring FALSE.
@@
Once the individual coding was completed, the two analysts performed an inter-coder agreement to obtain a common decision for each unit summary. Such inter-coder agreement allowed them to pinpoint discrepancies, but also find areas of improvement to fine-tune the LLM system. In the end, the inter-coder agreement was high, with an average of 95.05%. A merged database was created resulting from the agreement. This merged database was used to determine the performance, in terms of accuracy, precision and recall, of the LLM system in coding each of the CTI variables from the daily conversations on cybercrime forums. The results are presented below.
2.4 Accuracy, Precision and Recall
Accuracy measures the proportion of correct predictions (both true positives (TP) and true negatives (TN)) out of the total prediction [5]. It is used as a primary metric for predictive systems. However, in case of imbalanced datasets, where the majority class vastly outweighs the minority, accuracy can be misleading. A system might achieve high accuracy by focusing on the majority class while missing important minority class instances, which are often the most critical in cybersecurity. This is where precision and recall offer more insight into the system’s performance [5].
To give an example, take the variable is_remotely_exploitable sampled in a way that eight messages out of 500 are set to TRUE. Then, a system predicting FALSE 100% of the time would yield an accuracy of 492/500 = 98.4%, but the system is effectively useless as it can’t be used to detect remotely exploitable vulnerabilities as part of messages. Several solutions exist in order to better assess the true ability of the model to detect the needed signal.
Precision and Recall are two complementary metrics that offer a more nuanced understanding of the underlying phenomena, especially when working with imbalanced datasets. Precision is defined as the proportion of correctly identified true instances out of all instances the system predicted as positive. In contrast, recall refers to the system’s ability to capture all actual true cases from the dataset.
Precision is often used when predictions are costly or risky, and when the cost of ignoring false negatives (FN) predictions is lesser than the cost of predicting the false elements as being true. Recall, on the other hand, is often used when the cost of missing a prediction is very high, for example, in cancer detection or in detecting active cyber threats [5].
In cybersecurity, especially when dealing with real-life threat detection where the consequences of missing threats or mis-identifying signals can be extreme, these metrics are necessary to correctly understand the actual performance and reliability of the system. A system that prioritizes precision over recall may be suitable when false positives (FP) are highly undesirable, as could be the case when investigating threat actors in cybercrime forums. On the other hand, high recall is critical when it is more important not to miss any single threats, even at the cost of a potentially large number of false positives (FP), as in a critical infrastructure security operations environment.
In our context of CTI variable extraction, below is an interpretation of the performance of the LLM system tested. Note that since the original scope of the research was to identify geopolitical conversations in cybercrime forums, we first selected a balanced sample of the is_geopolitics variable and then assessed the system’s extraction capability for the other variables on this sample, leading to the varying levels of class imbalance shown below.
3 Assessing the Performance of the LLM System for CTI
The LLM system performed well in coding variables that represented valuable CTI information from cybercrime forums. Indeed, on average, the LLM was accurate in 96.23% of cases, with a minimum accuracy of 92.2% for the variable targeted_technologies and a maximum of 98.9% for the variable industries. For binary variables, the average precision was 90%, with a minimum of 75% and a maximum of 96.1%. The average recall was 88.2% with a minimum of 67.6% and a maximum of 99.2%. The percentage performance metrics for each variable are presented in Table 2 for the binary variables and in Table 3 for character variables.
Column N Precision Recall Accuracy TP TN FP FN
is_geopolitics 617 0.922 0.976 0.956 247 343 21 6
is_sale 702 0.909 0.992 0.962 250 425 25 2
is_initial_access 617 0.961 0.952 0.985 99 509 4 5
is_targeting_large_organizations 702 0.957 0.815 0.949 132 534 6 30
is_targeting_critical_infrastructure 617 0.852 0.676 0.927 69 503 12 33
is_remotely_exploitable 702 0.949 0.822 0.972 74 608 4 16
is_actively_exploited 702 0.750 0.941 0.973 48 635 16 3
Table 2: Performance Results for Binary Variables
@@ 3 Assessing the Performance of the LLM System for CTI
summary 500 0.988
targeted_technologies 231 0.922
industries 365 0.989
Table 3: Performance for Character Variables
For the summary variable, the coders ensured that the summary provided by the LLM system was accurate. The summary was accurate 98.8% (N=500) of the time. However, coders occasionally noticed missing key CTI information from the summaries. This is likely due to the inherent nature of summarization, which involves omitting some details. Also, what is considered essential can vary between coders.
The variable is_sale was coded well by the LLM system, achieving a precision of 90.9% and a recall of 99.2%. This means that the system accurately identified sales-related conversations with few false positives (precision), while nearly all true sales conversations were detected (high recall). Errors occurred when the LLM mistakenly flagged discussions about sales interest or general sales-related talk as actual sales. The variable is_initial_access performed similarly well, with a precision of 96.1% and recall of 95.2%, capturing almost all cases of initial access sales. The high precision and recall suggest that the LLM is reliable in detecting conversations around sales and access-related activities in cybercrime forums.
The variable is_targeting_large_organization had a precision of 95.7% but a lower recall of 81.5%, indicating that while most flagged conversations were correctly identified as targeting large organizations, the LLM missed some relevant conversations (lower recall). Similarly, the variable is_targeting_critical_infrastructure had a precision of 85.2% and a lower recall of 67.6%. The errors in these two variables often stemmed from difficulties in defining what qualifies as a “large organization” or a “critical infrastructure.” Despite these issues, analysts can still use these variables to get an overall sense of whether large organizations or critical infrastructures are being discussed in cybercrime forums.
The variable is_remotely_exploitable flagged conversations about remotely exploitable vulnerabilities, with a precision of 94.9% and recall of 82.2%. This suggests that most flagged vulnerabilities were relevant (high precision), but some vulnerabilities were missed (lower recall). On the other hand, the variable is_actively_exploited achieved a precision of 75% and recall of 94.1%, meaning that while some false positives were generated (lower precision), most conversations about actively exploited vulnerabilities were detected (high recall). Together, these two variables can help analysts identify vulnerabilities that organizations should prioritize for remediation.
The variable is_geopolitics flagged conversations related to geopolitical issues and was coded with a precision of 92.2% and recall of 97.6%. This high recall means that the LLM captured almost all relevant conversations. However, the precision was slightly lower because the LLM tended to flag the variable as TRUE whenever a country was mentioned, even when the discussion wasn’t strictly geopolitical. In this case, analysts can use this variable to focus on conversations that may impact specific countries or geopolitical conflicts.
@@
For the variable targeted_technologies, coders ensured that the technologies mentioned in messages matched those identified by the LLM. With an accuracy of 92.2%, the LLM rarely missed relevant technologies but occasionally flagged some incorrectly. Given this high accuracy, analysts can confidently use this variable to monitor which technologies are being discussed in cybercrime forums.
Finally, the variable industries achieved an accuracy of 98.9%, meaning that almost every industry mention was correctly identified without any false positives (FP) or false negatives (FN). This high score might be due to the better-defined nature of this variable, contributing to the LLM’s effectiveness in identifying industry mentions. Analysts and policymakers can use this variable to track discussions in cybercrime forums that target specific industries, facilitating better risk assessments.
In conclusion, the differences between precision and recall emphasize the interpretative nature of the variables. Precision reflects how accurately the LLM identifies relevant cases without false positives (FP), while recall shows its ability to capture all relevant instances. High precision and recall for variables like is_sale and is_initial_access indicate reliable performance. However, variables like is_targeting_critical_infrastructure have lower recall, meaning relevant conversations were sometimes missed, highlighting the need for improvement in capturing more nuanced discussions, particularly around complex definitions like “critical infrastructure.”
@@ 4 Coders’ Insights to Fine-Tune the LLM System
Given the results of the analysis, there is no doubt that LLM systems can be useful in extracting key CTI information from cybercrime forums. Indeed, the summaries generated from daily conversations focused on relevant information, even when a large number of messages were posted. Such relevant information was sometimes even missed by the analysts.
Reviewing the same conversations allowed the coders to discuss and identify areas for improvement in coding CTI variables. These areas are useful for any researcher or analyst wishing to use LLMs to code variables based on text input. They are presented below.
@@ 4.2 Verb Tenses in Prompts
Most likely, the verb tense in the prompt, “an actor is selling something,” was the reason why the LLM did not code this conversation as involving a sale. However, upon reflection, it could be considered a sale because the item in question has been sold. In the end, if the goal is to code all sales, past or present, it might be necessary to revisit the verb tenses in the prompts given to the LLM.
@@ 4.5 Links between Variables with Similar Concepts
In this example, the LLM coded FALSE for is_targeting_critical_infrastructure and TRUE for is_targeting_large_organization. In the industries variable, the LLM indicated [’Critical Infrastructure’, ’Military’]. This discrepancy raises questions about the definitions of these three variables, as discussed above, and how the LLM interpreted them. It also showcases that some variables may be used to code others. For example, a technology mentioned in the summary was not always associated with an organization, despite the obvious connection between the two. In this excerpt:
@@ 4.7 The LLM System is Imperfect, just like Humans
The LLM coded the variable is_initial_access as FALSE even though the actor sells a tool that allows remote control over a target machine via a web browser. Hence, in some cases, the LLM missed the information. The LLM system was therefore not flawless. Although rare, it made coding errors that could not be explained.